A total charge of 6.3×10−8 C is distributed uniformly throughout a 2.7-cm radius sphere. What is the volume charge density?
1
Expert's answer
2020-03-31T09:16:07-0400

"q=6.3\times10^{-8} C"

"r=2.7" cm "=2.7\times10^{-2} m"

let's first find volume "V={\frac 4 3}\pi r^3=8.24\times10^{-5} m^3"

Then volume charge density "\rho=q/V=7.65\times10^{-4} {\frac C {m^3}}"